# ReefNet: A Large-scale, Taxonomically Enriched Dataset and Benchmark for Hard Coral Classification

### v1

@@ A.9 Resources Availibity
CREP-REA HAWAII NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018b) Annette DesRochers, Andrew Gray, Brett Schumacher, Bernardo Vargas Angel, Courtney Couch, Ivor Williams, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Tom Oliver, Troy Kanemura, Ari Halperin, Chelsie Counsell, Colt Davis, Isabelle Basden, Jon Ehrenberg, Kerry Reardon, Mia Lamirand, Roseanna Lee National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
CREP-REA MARIANAS NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018a) Annette DesRochers, Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Tom Oliver, Troy Kanemura, Kaylyn McCoy, Kevin Lino, Marie Ferguson, Mia Lamirand, Nalani Kito-Ho, Winter Jimenez, Brett Schumacher National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
CREP-REA NWHI 2017 Tiny Photos NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018b) Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jon Ehrenberg National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
CREP-REA SAMOA/PRIA NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018c) Annette DesRochers, Andrew Gray, Brett Schumacher, Bernardo Vargas Angel, Courtney Couch, Ivor Williams, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Tom Oliver, Troy Kanemura, Isabelle Basden, Mia Lamirand, Andrew Shantz, Brittany Huntington, Corinne Amir, Hatsue Bailey, Marie Ferguson, Mollie Asbury, Nalani Kito-Ho, Winter Jiminez, Helen Ford, Nicole Kamalu National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
@@
ESD_REA HAWAII_v2 NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018b) Annette DesRochers, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Tom Oliver, Isabelle Basden, Jon Ehrenberg, Hatsue Bailey, John Morris, Mia Larimand, Paula Misa National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
ESD_REA_MARIANA_v2 NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018a) Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Kaylyn McCoy, Tom Oliver, Ari Halperin, Mia Lamirand, Hatsue Bailey, Nicolas Osborn, Jon Ehrenberg National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
ESD_REA_SAMOA_PRIA_v2 NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018c) Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Ari Halperin, Isabelle Basden, Mia Lamirand, Hatsue Bailey, Nicolas Osborn, Tom Oliver National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
@@
NOAA ESD Coral Bleaching Classifier  Ehrenberg et al. (2022) Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Tom Oliver, Jon Ehrenberg National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
@@
Tonga_2022-08 & Samoa_2022-09 & Samoa_2019-12 & Samoa_2017-12NOAA Pacific Islands Fisheries Science Center, Ecosystem Sciences Division (2018c) Chris Sullivan, Katie Lubarsky, Gloria Mariño-Briceño, Hannah Gower, Kylie Yogi, Phi Lang Scripps Institution of Oceanography, University of California San Diego

### v2

@@ A.9 Resources Availibity
CREP-REA HAWAII NOAA_NCRMP_HI_2018 Annette DesRochers, Andrew Gray, Brett Schumacher, Bernardo Vargas Angel, Courtney Couch, Ivor Williams, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Tom Oliver, Troy Kanemura, Ari Halperin, Chelsie Counsell, Colt Davis, Isabelle Basden, Jon Ehrenberg, Kerry Reardon, Mia Lamirand, Roseanna Lee National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
CREP-REA MARIANAS NOAA_NCRMP_2018 Annette DesRochers, Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Tom Oliver, Troy Kanemura, Kaylyn McCoy, Kevin Lino, Marie Ferguson, Mia Lamirand, Nalani Kito-Ho, Winter Jimenez, Brett Schumacher National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
CREP-REA NWHI 2017 Tiny Photos NOAA_NCRMP_HI_2018 Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jon Ehrenberg National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
CREP-REA SAMOA/PRIA NOAA_NCRMP_SAMOA_2018 Annette DesRochers, Andrew Gray, Brett Schumacher, Bernardo Vargas Angel, Courtney Couch, Ivor Williams, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Tom Oliver, Troy Kanemura, Isabelle Basden, Mia Lamirand, Andrew Shantz, Brittany Huntington, Corinne Amir, Hatsue Bailey, Marie Ferguson, Mollie Asbury, Nalani Kito-Ho, Winter Jiminez, Helen Ford, Nicole Kamalu National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
@@
ESD_REA HAWAII_v2 NOAA_NCRMP_HI_2018 Annette DesRochers, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Tom Oliver, Isabelle Basden, Jon Ehrenberg, Hatsue Bailey, John Morris, Mia Larimand, Paula Misa National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
ESD_REA_MARIANA_v2 NOAA_NCRMP_2018 Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Kaylyn McCoy, Tom Oliver, Ari Halperin, Mia Lamirand, Hatsue Bailey, Nicolas Osborn, Jon Ehrenberg National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
ESD_REA_SAMOA_PRIA_v2 NOAA_NCRMP_SAMOA_2018 Andrew Gray, Bernardo Vargas Angel, Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Paula Misa, Ari Halperin, Isabelle Basden, Mia Lamirand, Hatsue Bailey, Nicolas Osborn, Tom Oliver National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
@@
NOAA ESD Coral Bleaching Classifier  Ehrenberg2022 Courtney Couch, Jonathan Charendoff, Morgan Winston Pomeroy, Tom Oliver, Jon Ehrenberg National Oceanographic and Atmospheric Administration, Pacific Islands Fisheries Science Center
@@
Tonga_2022-08 & Samoa_2022-09 & Samoa_2019-12 & Samoa_2017-12NOAA_NCRMP_SAMOA_2018 Chris Sullivan, Katie Lubarsky, Gloria Mariño-Briceño, Hannah Gower, Kylie Yogi, Phi Lang Scripps Institution of Oceanography, University of California San Diego